OpenStudy (anonymous):

okay so i wrote out a table for h(x) and the function coordinates (-4,0), (-2,3), (0,0), (2,5), (4,2), and (6,2)...now i have to translate this function to 2h(x-5)+3...i know i have to add 5 to the x's in the table but what do i change with the y's in the table? do i add 2 since there's a 2 infront of the h in the function? or do i add 3 since it says +3?

OpenStudy (anonymous):

the thing is like this y =h(x) and point (-4,0) means h(-4)=0=y now h(x-5)=h(-4-5=h(-9) u have to solve this way

OpenStudy (anonymous):

but like in general when there is a number before h and i have a table of coordinates, do i multply the y's by the coefficent of h?

OpenStudy (anonymous):

or do i add the "k" or 3 in this function

OpenStudy (anonymous):

first multiply the coefficient then add or substract

OpenStudy (anonymous):

okay so i multiply 2 to all of the y's in the table and then add 3?

OpenStudy (anonymous):

\[2h(x-5)+3\] translates by shifting to the right 5, doubling and then adding 3

OpenStudy (anonymous):

yeah but what do i do to the table for the y's??

OpenStudy (anonymous):

??